OpenStudy (anonymous):

Use the given zero to find remaining zeros. Show work. h(x)=x^4-12x^3+28x^2+196x-1845 zero: 4-5i

OpenStudy (anonymous):

Okay ill give it a stap.. So they told you that 4-5i was a root, the funny thing about complex/imaginary roots they come as pairs. More formally called conjugate pairs. I claim that an other root is 4+5i since it will complete the pair. \[x^4-12x^3+28x^2+196x-1845=(x-4-5i)(x+4-5i)(ax^2+bx+c)\] \[= ax^4+(b-8a)x^3+(41a-8b+c)x^2+(41b-8c)x+41c\] Now set the coefficients of the orginal equation to the new one interms of a,b,c. \[a=1\] \[b-8a= -12\] \[41a-8b+c=28\] \[41b-8c=196\] \[41c =- 1845\] Solve for a,b,c and then all you need to do is find the roots of \[ax^2+bx+c.\] which will be \[x=-5 \]and \[x= 9\]